### v1

@@ 5 Conclusion
In this paper, we investigated quantum corrections to the mean-field dynamics of the relative phase between two coupled condensates. Starting from the classical action expressed in the collective variables $(\phi,z)$, we obtained an effective action depending only on the phase, and we derived the corresponding one-loop quantum effective action by means of a covariant background-field method that fully accounts for the spatial dependence of the mass and the exact form of the potential. Comparison with the exact dynamics of the two-site Bose-Hubbard model shows that the one-loop corrected dynamics reproduces the quantum evolution significantly better than the standard mean-field equations over a wide region of the parameter space $(U/J,N)$. For small oscillations around the equilibrium value $\phi=0$, the dynamics is harmonic, as in the mean-field case, but with a modified Josephson frequency. The shift of the classical frequency induced by quantum fluctuations lies between $1\%$ and $3\%$ for the parameter range considered in Fig. 3.
Besides elucidating the correspondence between the quantum theory and its classical limit, and quantifying the importance of quantum effects for different interaction strengths and number of particles, the corrected Josephson frequency may be employed to enhance the accuracy of semiclassical schemes used in the description of thermodynamic properties of these systems [25]. Future work could generalize the quantum effective-action approach to the full set of collective variables $(\phi,z)$, thereby allowing a systematic study of quantum corrections to phenomena such as macroscopic quantum self-trapping, and could extend the present treatment to finite temperature.

### v2

@@ 5 Conclusion
In this paper, we investigated quantum corrections to the mean-field dynamics of the relative phase between two coupled condensates. Starting from the classical action expressed in the collective variables $(\phi,z)$, we obtained an effective action depending only on the phase, and we derived the corresponding one-loop quantum effective action by means of a covariant background-field method that fully accounts for the spatial dependence of the mass and the exact form of the potential. Comparison with the exact dynamics of the two-site Bose-Hubbard model shows that the one-loop corrected dynamics reproduces the quantum evolution significantly better than the standard mean-field equations over a wide region of the parameter space (N,U/J)(N,\,U/J), confirming that the field-theoretic quantum effective action approach provides a systematic and physically transparent framework to capture leading-order quantum effects in Josephson junctions. For small oscillations around the equilibrium value zero, the dynamics is harmonic, as in the mean-field case, but with a modified Josephson frequency. The shift of the classical frequency induced by quantum fluctuations lies between $1\%$ and $3\%$ for the parameter range considered in Fig. 3.
Besides elucidating the correspondence between the quantum theory and its classical limit, and quantifying the importance of quantum effects for different interaction strengths and number of particles, the corrected Josephson frequency may be employed to enhance the accuracy of semiclassical schemes that can be used in the description of thermodynamic properties of these systems [25]. Future work could generalize the quantum effective-action approach to the full set of collective variables $(\phi,z)$, thereby allowing a systematic study of quantum corrections to phenomena such as macroscopic quantum self-trapping, and could extend the present treatment to finite temperature.